# Algorithms for Reconstructing B Cell Lineages in the Presence of Context-Dependent Somatic Hypermutation

### v1

@@ 5.2 Effects of context-dependence on tree and UCA reconstruction accuracy
Improving UCA reconstruction with VDJ recombination models: In standard phylogenetics, the common ancestor $\mathbf{x}$ is assumed to be drawn from the stationary distribution of the evolutionary model. In affinity maturation, the UCA is generated by the mechanisms of VDJ recombination. [hozumi_evidence_1976, tonegawa_somatic_1983]. Detailed stochastic models of this process have been developed [dhar_bayesian_2020, kepler_reconstructing_2013, Marcou2018] that assign a probability to every possible UCA sequence arising from the host germline VDJ gene segment library via recombination and $n$-nucleotide insertion. Although developed for BCR sequence analysis including the specific purpose of lineage reconstruction and UCA estimation, these models have relied primarily on independent-site models of SHM. However, VDJ models can provide additional information regarding the UCA prior to SHM, and incorporating this information may be expected to improve reconstruction accuracy. Table 3a shows the results of incorporating the HMM-based VDJ rearrangement model of [dhar_bayesian_2020]. As expected, the accuracy of UCA reconstruction does indeed significantly improve, although accuracy obtained by our importance sampling-based DSM estimator is constrained by the failure of the ISM-based linearham model to sample the correct UCA at all. Notably, all observed reconstruction errors occur in the CDR3 region, suggesting that the prior information of the VDJ model significantly improves accuracy in the germline-templated V, and J regions, but is less accurate in the highly variable $n$-nucleotide inserted regions, where it may actually introduce errors. Piecewise VDJ prior: Table 3a shows that substituting the piecewise prior described in section 4.6 for the full VDJ prior does indeed provide a small further improvement in UCA reconstruction accuracy.
@@
VDJ Reparameterization: Results in Table 3a incorporating the VDJ prior were obtained using default parameters for linearham. To test the sensitivity of these results to the parameterization of the VDJ model, we re-ran simulation studies using a parameter set for linearham obtained by training on a large NGS dataset of human B cell repertoire heavy chain sequences. 10 independent simulated clonal datasets were generated using UCAs sampled from the stochastic VDJ rearrangement process (via linearham) and a tree topology (including branch lengths) sampled from the Yule prior. Observed sequences were then obtained by simulating context-dependent evolution under the ARMADiLLO model along the sampled tree topology. These simulated repertoires match many key statistics of real-world clones, with a typical clone having sequence length around 365 and containing 55 highly-mutated observed sequences, with average Hamming distance from the UCA of 84.44 ($\approx 75\%$ sequence identity), average tree depth similar to that of trees reconstructed for the CH235 clone, and depth of the leaf closest to the root (which provides most information about the UCA) of $0.08$ with 27.35 mutations from the root sequence on average (least mutated sequence in real CH235 clone dataset: 27 mutations). SMC was used in all reconstructions based on results in Table 2.
@@
Observed ISM mistakes To understand the impact of the independent site assumption in the presence of site-dependence, we identified mistakes commonly observed under ISM reconstruction. These illustrate how failure to account for context-dependent mutation leads to inaccuracies in phylogenetic tree reconstruction. Scenario 1: Figure 5(a) shows a small subtree of a larger tree, with $\mathrm{S}_{1560}$, $\mathrm{S}_{1561}$ and $\mathrm{S}_{4211}$ observed leaves. It thus represents a tree reconstruction problem with only 3 leaf nodes and observed sequences of length 367. Despite the large amount of information, the ISM-reconstructed topology is incorrect. Under the ISM, the reconstructed tree minimizes the sum of Hamming distances along edges. Key sites are 147 and 188: Figure 5 shows that $\mathrm{S}_{1561}$ and $\mathrm{S}_{4211}$ agree on $\mathrm{T}$ at site $146$ while $\mathrm{S}_{1560}$ and $\mathrm{S}_{4211}$ agree on $\mathrm{C}$ at site $188$, hence the ISM estimates the UCA as $\mathrm{T}$ at site $147$ and $\mathrm{C}$ at site $188$ with high probability, ensuring each site has only experienced a single mutation. However the DSM assigns a lower probability to the ISM tree since these two key mutations occur in highly mutable motifs. As seen in Figure 5(b), the ISM tree requires site $147$ to remain unmutated along the path $\mathrm{E}\rightarrow\mathrm{F}\rightarrow\mathrm{S}_{1561}$ and the path $\mathrm{E}\rightarrow\mathrm{F}\rightarrow\mathrm{S}_{4211}$, which is far less likely under the DSM. Scenario 2: In the example of Figure 5(d), the (sub)tree obtained under the ISM has Hamming distance 8 between the UCA (node $\mathrm{A}$) and $\mathrm{S}_{80}$. Since 4 of those 8 mutations are shared with $\mathrm{S}_{245}$, the ISM places $\mathrm{S}_{80}$ and $\mathrm{S}_{245}$ on the same lineage $A\rightarrow D$. However, the contexts of these 4 mutations are highly mutable, and under the DSM they are more likely to have been duplicated independently along separate lineages.
@@ 5.3 Analysis of HIV bnAb B cell repertoires
DH270: The UCA estimated for the DH270 clone under the DSM differs by only a single amino acid from the UCA estimated under the ISM, in 7th position of CDRH3 (Fig. 6a, red). This site, which is reconstructed as a glycine under the ISM and a serine under the context-dependent model, occurs unambiguously in the non-templated nucleotide (n-nucleotide) regions that are not encoded by the V, D or J gene segments and thus cannot be resolved by priors on the initial VDJ rearrangement. While this difference in reconstructions would appear to be small, in fact the 7th position in the CRH3 is critical for binding of DH270 lineage members to the HIV Env protein as it makes structural contacts to the V3 glycan epitope [Henderson2023]. Moreover, deep scanning mutagenesis in the CDRH3 of a previously inferred DH270 UCA revealed that the 7th position in the CDRH3 could only tolerate a serine to maintain binding to the designed priming immunogen  [swanson_identification_2023]. Thus, the one position difference between the ISM and DSM reconstructions occurs at a key site for both bnAb epitope recognition and for binding to an immunogen designed to activate the DH270 lineage, and represents a critical error of the ISM assumption. This example underscores the importance of highly accurate UCA estimates for both understanding the natural maturation history of clonal lineages and for vaccine design. While the true UCA sequence for the DH270 clone is unknowable, the fact that the DSM-estimated UCA contains the serine at this critical 7th position which is the only amino acid compatible with binding to the designed priming immunogen – itself a modified form of an HIV Env circulating at the time the DH270 clone emerged – suggests that the DSM-estimated UCA may more accurately reflect the true ancestral sequence. Moreover, designing an immunogen to elicit this bnAb based on the incorrect ISM-reconstructed UCA could be counterproductive, potentially selecting for mutants that compete against the desired bnAb.
@@
CH235 We observe two amino acid differences between the ISM and DSM reconstructions of the CH235 UCA CDRH3, at positions 99 and 102 (Fig. 6b). The CH235 CDRH3 forms extensive structural contacts with the epitope and is critical for recognition of the conserved CD4 binding site [Henderson_2024]. Position 102 makes direct contact with the CH235 epitope, and position 99 likely influences CDRH3 positioning which may be critical for epitope recognition. This underscores the importance of accurate inference of the unmutated CDRH3 in order to identify Env variants capable of engaging CH235 precursor B cells, a key step in immunogen design.
@@ Appendix 0.F Experiments
For the DSM, we consistently use the mutation rates specified by the ARMADiLLO model [wiehe_functional_2018]. For the ISM, we employ the corresponding mean-field approximation. This choice helps reduce the possibility that differences in the inferred trees or UCAs between the ISM and DSM arise from disparities in the underlying rate parameters rather than from the presence or absence of context-dependent effects.

### v2

@@ 5.2.3 Improving UCA reconstruction with VDJ recombination models:
In standard phylogenetics, the common ancestor $\mathbf{x}$ is assumed to be drawn from the stationary distribution of the evolutionary model. In affinity maturation, the UCA is generated by the mechanisms of VDJ recombination. [hozumi_evidence_1976, tonegawa_somatic_1983]. Detailed stochastic models of this process have been developed [dhar_bayesian_2020, kepler_reconstructing_2013, Marcou2018] that assign a probability to every possible UCA sequence arising from the host germline VDJ gene segment library via recombination and $n$-nucleotide insertion. Although developed for BCR sequence analysis including the specific purpose of lineage reconstruction and UCA estimation, these models have relied primarily on independent-site models of SHM. However, VDJ models can provide additional information regarding the UCA prior to SHM, and incorporating this information may be expected to improve reconstruction accuracy. Table 3a shows the results of incorporating the HMM-based VDJ rearrangement model of [dhar_bayesian_2020]. As expected, the accuracy of UCA reconstruction does indeed significantly improve, although accuracy obtained by our importance sampling-based DSM estimator is constrained by the failure of the ISM-based linearham model to sample the correct UCA at all. Notably, all observed reconstruction errors occur in the CDR3 region, suggesting that the prior information of the VDJ model significantly improves accuracy in the germline-templated V, and J regions, but is less accurate in the highly variable $n$-nucleotide inserted regions, where it may actually introduce errors. Piecewise VDJ prior: Table 3a shows that substituting the piecewise prior described in section 4.6 for the full VDJ prior does indeed provide a small further improvement in UCA reconstruction accuracy.
@@
5.2.4 VDJ Reparameterization:
Results in Table 3a incorporating the VDJ prior were obtained using default parameters for linearham. To test the sensitivity of these results to the parameterization of the VDJ model, we re-ran simulation studies using a parameter set for linearham obtained by training on a large NGS dataset of human B cell repertoire heavy chain sequences. 10 independent simulated clonal datasets were generated using UCAs sampled from the stochastic VDJ rearrangement process (via linearham) and a tree topology (including branch lengths) sampled from the Yule prior. Observed sequences were then obtained by simulating context-dependent evolution under the ARMADiLLO model along the sampled tree topology. These simulated repertoires match many key statistics of real-world clones, with a typical clone having sequence length around 365 and containing 55 highly-mutated observed sequences, with average Hamming distance from the UCA of 84.44 ($\approx 75\%$ sequence identity), average tree depth similar to that of trees reconstructed for the CH235 clone, and depth of the leaf closest to the root (which provides most information about the UCA) of $0.08$ with 27.35 mutations from the root sequence on average (least mutated sequence in real CH235 clone dataset: 27 mutations). SMC was used in all reconstructions based on results in Table 2.
@@ 5.2.4 VDJ Reparameterization:
5.2.5 Observed ISM mistakes
To understand the impact of the independent site assumption in the presence of site-dependence, we identified mistakes commonly observed under ISM reconstruction. These illustrate how failure to account for context-dependent mutation leads to inaccuracies in phylogenetic tree reconstruction.
Scenario 1:
Figure 5(a) shows a small subtree of a larger tree, with $\mathrm{S}_{1560}$, $\mathrm{S}_{1561}$ and $\mathrm{S}_{4211}$ observed leaves. It thus represents a tree reconstruction problem with only 3 leaf nodes and observed sequences of length 367. Despite the large amount of information, the ISM-reconstructed topology is incorrect. Under the ISM, the reconstructed tree minimizes the sum of Hamming distances along edges. Key sites are 147 and 188: Figure 5 shows that $\mathrm{S}_{1561}$ and $\mathrm{S}_{4211}$ agree on $\mathrm{T}$ at site $146$ while $\mathrm{S}_{1560}$ and $\mathrm{S}_{4211}$ agree on $\mathrm{C}$ at site $188$, hence the ISM estimates the UCA as $\mathrm{T}$ at site $147$ and $\mathrm{C}$ at site $188$ with high probability, ensuring each site has only experienced a single mutation. However the DSM assigns a lower probability to the ISM tree since these two key mutations occur in highly mutable motifs. As seen in Figure 5(b), the ISM tree requires site $147$ to remain unmutated along the path $\mathrm{E}\rightarrow\mathrm{F}\rightarrow\mathrm{S}_{1561}$ and the path $\mathrm{E}\rightarrow\mathrm{F}\rightarrow\mathrm{S}_{4211}$, which is far less likely under the DSM.
Scenario 2:
In the example of Figure 5(d), the (sub)tree obtained under the ISM has Hamming distance 8 between the UCA (node $\mathrm{A}$) and $\mathrm{S}_{80}$. Since 4 of those 8 mutations are shared with $\mathrm{S}_{245}$, the ISM places $\mathrm{S}_{80}$ and $\mathrm{S}_{245}$ on the same lineage $A\rightarrow D$. However, the contexts of these 4 mutations are highly mutable, and under the DSM they are more likely to have been duplicated independently along separate lineages.
@@ 5.3 Analysis of HIV bnAb B cell repertoires
5.3.1 DH270
The UCA estimated for the DH270 clone under the DSM differs by only a single amino acid from the UCA estimated under the ISM, in 7th position of CDRH3 (Fig. 6a, red). This site, which is reconstructed as a glycine under the ISM and a serine under the context-dependent model, occurs unambiguously in the non-templated nucleotide (n-nucleotide) regions that are not encoded by the V, D or J gene segments and thus cannot be resolved by priors on the initial VDJ rearrangement. While this difference in reconstructions would appear to be small, in fact the 7th position in the CRH3 is critical for binding of DH270 lineage members to the HIV Env protein as it makes structural contacts to the V3 glycan epitope [Henderson2023]. Moreover, deep scanning mutagenesis in the CDRH3 of a previously inferred DH270 UCA revealed that the 7th position in the CDRH3 could only tolerate a serine to maintain binding to the designed priming immunogen  [swanson_identification_2023]. Thus, the one position difference between the ISM and DSM reconstructions occurs at a key site for both bnAb epitope recognition and for binding to an immunogen designed to activate the DH270 lineage, and represents a critical error of the ISM assumption. This example underscores the importance of highly accurate UCA estimates for both understanding the natural maturation history of clonal lineages and for vaccine design. While the true UCA sequence for the DH270 clone is unknowable, the fact that the DSM-estimated UCA contains the serine at this critical 7th position which is the only amino acid compatible with binding to the designed priming immunogen – itself a modified form of an HIV Env circulating at the time the DH270 clone emerged – suggests that the DSM-estimated UCA may more accurately reflect the true ancestral sequence. Moreover, designing an immunogen to elicit this bnAb based on the incorrect ISM-reconstructed UCA could be counterproductive, potentially selecting for mutants that compete against the desired bnAb.
@@ 5.3.1 DH270
5.3.2 CH235
We observe two amino acid differences between the ISM and DSM reconstructions of the CH235 UCA CDRH3, at positions 99 and 102 (Fig. 6b). The CH235 CDRH3 forms extensive structural contacts with the epitope and is critical for recognition of the conserved CD4 binding site [Henderson_2024]. Position 102 makes direct contact with the CH235 epitope, and position 99 likely influences CDRH3 positioning which may be critical for epitope recognition. This underscores the importance of accurate inference of the unmutated CDRH3 in order to identify Env variants capable of engaging CH235 precursor B cells, a key step in immunogen design.
@@ Appendix 0.F Experiments
For the DSM, we consistently use the mutation rates specified by the ARMADiLLO model [Wiehe:2018]. For the ISM, we employ the corresponding mean-field approximation. This choice helps reduce the possibility that differences in the inferred trees or UCAs between the ISM and DSM arise from disparities in the underlying rate parameters rather than from the presence or absence of context-dependent effects.